# Gibbs polystability of Fano manifolds, stability thresholds and symmetry breaking

### v2

@@ 4.1. Spontaneous symmetry breaking
although both $F_{\gamma}$ and $\mathcal{P}(X)_{0}$ are $S^{1}$–invariant (and likewise for $\mathcal{D}_{\gamma}$ and $C^{1}(X)_{0})$. The point is that the destabilizing curve $\mu^{t}$, exhibited in the proof of Theorem 3.13, is not $S^{1}$–invariant (with respect to the $S^{1}$–action fixing 0 and $\infty).$ This is in contrast to the case when $\gamma=1,$ where the two infima above do coincide. Indeed, when $\gamma=-1$, the inf of $F_{\gamma}$ is attained at the unique Kähler–Einstein volume form in $\mathcal{P}(X)_{0}$ and this volume form is $S^{1}$–invariant (and likewise for $\mathcal{D}_{1}(u)).$
@@ 4.1.1. Failure of symmetric rearrangements
In [37, 38] and [9], the sharp logarithmic HLS inequality on the two-sphere $X$ (corresponding to the case $w=0$ and $\gamma=1)$ was shown using symmetric decreasing rearrangements. The arrangement was defined with respect to Lebesgue measure on $\mathbb{R}^{2}$ in [37, 38] (using stereographic projection), while in [9] it was defined with respect to the $\mathrm{SO}(3)$–invariant measure on the two-sphere. Both these two rearrangements define maps $\mathcal{P}(X)\rightarrow\mathcal{P}(X)^{S^{1}}$ decreasing the functional $F_{1}$ (and the arrangement in [9] decreases $F_{\gamma}$ for any $\gamma>0).$ However, neither of the rearrangements preserve the space $\mathcal{P}(X)_{0}$ (for any given weight $w).$ Moreover, the spontaneous symmetry breaking exhibited in formula 4.11 implies that when $w\in]0,1/2[$ and $\gamma\in]1/(1-w),2[$ there is no map $\mathcal{P}(X)_{0}\rightarrow\mathcal{P}(X)_{0}^{S^{1}}$ that decreases the functional $F_{\gamma}.$
@@ 4.2. Restricting to $\mathcal{K}$-invariant metrics
By [77, Thm 2.18], $\delta_{C(\mathcal{G})}^{A}(X,\Delta)>1$ iff $(X,\Delta)$ admits a KE metric. Hence, by [48, Thm 2.18], $\delta_{C(\mathcal{G})}^{A}(X,\Delta)>1$ iff $\delta_{\mathcal{G}}^{A}(X,\Delta)>1.$ However, in general these two analytic invariants may differ:
@@
5. A sharp constrained log HLS inequality on $S^{2}$
In this section we will prove the following result, which is equivalent to Theorem 1.4 stated in the introduction (by formula 3.13 and Prop 3.4):
@@ 5. A sharp constrained log HLS inequality on $S^{2}$
Let $X$ be the two-sphere, $\Delta=0$ and assume that $\phi_{0}$ is $\mathrm{SO}(3)$–invariant. Then for any given $\gamma\in]0,2]$
$F_{\gamma}(\mu)\geq 0\,\,\,\forall\mu\in\mathcal{P}(X)_{0},$
[…]
@@
When $\gamma=1$ the inequality for $F_{1}(\mu)$ follows from the general fact that, when a Fano manifold $X$ admits a Kähler–Einstein metric $\omega_{\text{KE}}\in C^{1}(X),$ the Mabuchi functional $\mathcal{M}$ is minimized on the Kähler potential $u$ of $\omega_{\text{KE}}$ [8] and $\mathcal{M}$ is then $\mathcal{G}$-invariant – so that the moment condition on $\mu$ can be dispensed with. In the case of the two-sphere (or more generally, the unit sphere in $\mathbb{R}^{n+1})$ the inequality $F_{1}(\mu)\geq 0$ on $\mathcal{P}(X)$ is shown in [37, 38, 9] using symmetrization (the inequality $F_{\gamma}(\mu)\geq 0$ for $\gamma\in[0,1]$ then follows directly).
@@
Before embarking on the detailed proof of Theorem 5.1 it may be illuminating to compare its structure to the proof of Theorem 4.5 in [72]. The starting point of the latter is Aubin’s constrained Moser–Trudinger inequality which, in our notations, implies that the twisted Ding functional $\mathcal{D}_{\gamma}$ admits a minimizer $u_{\gamma}$ on the space of all functions in the Hilbert space $W^{2,1}(X)$ satisfying Aubin’s constraint. Since this is an infinite dimensional Hilbert manifold it follows that $u_{\gamma}$ satisfies an Euler–Lagrange equation, involving the differential of $\mathcal{D}_{\gamma}$ and three Lagrange multipliers (incorporating Aubin’s constraint). As shown in [39, Lemma 3.1], the Lagrange multipliers vanish and, as a result, $u_{\gamma}$ is, in fact, a critical point of $\mathcal{D}_{\gamma}$ on $W^{2,1}(X)$ and thus satisfies a Liouville equation on $S^{2},$ that is shown to have only have constant solutions in [72] (thus confirming a conjecture in [39]).
@@
In order to prove Theorem 5.1 we replace $\mathcal{D}_{\gamma}$ with the free energy functional $F_{\gamma}$ on $\mathcal{P}(X)_{0}.$ It follows from the formula $\delta^{\mathrm{A}}(X,0)^{\mathcal{G}}=2$ in Theorem 3.13 that $F_{\gamma}$ admits a minimizer $\mu_{\gamma}$ in $\mathcal{P}(X)_{0}$ for any given $\in]0,2[.$ However, since the space $\mathcal{P}(X)$ has a boundary it seems hard to directly show that $\mu_{\gamma}$ satisfies an appropriate equation. Instead, we leverage a duality formula, expressing the infimum of $F_{\gamma}$ over $\mathcal{P}(X)_{0}$ as the infimum over $W^{2,1}(X)$ of a perturbation of the twisted Ding functional:
$\inf_{\mathcal{P}(X)_{0}}F_{\gamma}=\gamma\inf_{u\in W^{2,1}(X)}\left(\mathcal{D}_{\gamma}(u)+\delta_{\gamma}(u)\right),$
@@
where $\delta_{\gamma}$ is an entropic distance, defining a function on $W^{2,1}(X)$ taking values in $0,\infty[,$ that vanishes precisely when $\gamma u$ satisfies Aubin’s constraint 4.8 (in particular, the unboundedness from above of $\delta_{\gamma}(u)$ compensates for the unboundedness from below of $\mathcal{D}_{\gamma}(u)).$ This bypasses the boundary problem and allows us to reduce the proof of Theorem 5.1 to the uniqueness result in [72]. In fact, the proof will show that the inf in the right hand side is attained precisely when $u$ is constant and thus vanishes.
@@
A consequence of the proof is that the infimum of $\mathcal{D}_{\gamma}(u)+\delta_{\gamma}(u)$ over $W^{2,1}(X)$ coincides with the one obtained by replacing the functional $\delta_{\gamma}(u)$ with the indicator function $\chi(u)$ attached to Aubin’s constraint (i.e. $\chi(u)$ vanishes when $\gamma u$ satisfies Aubin’s constraint and $\chi(u)=\infty$ otherwise). From a general optimization perspective this is somewhat miraculous.
@@
which defines a probability measure on $X,$ by the previous lemma. It will be enough to prove the inequality in Theorem 5.1 for a fixed $\gamma\in]1,2[.$ By Theorem 3.13 the infimum of $F_{\gamma}$ on $\mathcal{P}(X)_{0}$ is then finite. By the previous lemma we thus have
$\inf_{\mathcal{P}(X)_{0}}F_{\gamma}=\inf_{\mathcal{P}(X)_{0}\cap W^{2,1*}}F_{\gamma}>-\infty.$
@@ 5.1.1. A duality formula
(5.1) $\inf_{\mathcal{P}(X)_{0}\cap W^{2,1*}}F_{\gamma}=\gamma\inf_{u\in W^{2,1}(X)}\left(\mathcal{D}_{\gamma}(u)+\delta_{\gamma}(u)\right),\,\,\,\delta_{\gamma}(u)\coloneqq\frac{1}{\gamma}\inf_{\mu\in\mathcal{P}(X)_{0}}D(\mu|\mu_{\gamma u}),$
@@
then proves 5.4.
Next, note that the functional $\tilde{D}(\mu)$ on $W^{2,1}(X)^{*}$ is convex and lsc (since $\mathcal{P}(X)_{0}$ is compact and convex in $W^{2,1}(X)^{*})$ and $D(\mu|\mu_{0})$ is convex on $\mathcal{P}(X)_{0}).$ It thus follows from combining formula 5.4 with the duality 5.2 that
@@
so that formula 5.3 translates into
$\inf_{\mathcal{P}(X)_{0}\cap W^{2,1}(X)^{*}}F_{\gamma}=\inf_{W^{2,1}(X)^{*}}\left(-\mathcal{E}^{*}+\mathcal{A}^{*}\right).$
@@
By 5.2, this means that $\mathcal{A}(u)\geq\mathcal{E}(u)+C$ for any $u\in H.$ Conversely, if the latter inequality holds, then, by taking the sup over all $u\in H$ we directly deduce that $\mathcal{A}^{*}(\mu)\geq\mathcal{E^{*}}(\mu)+C$ for all $\mu\in H^{*}.$ Thus $(-\mathcal{E}^{*}+\mathcal{A}^{*})(\mu)\geq C$ for all $\mu\in H^{*}.$
In the present case $\mathcal{A}$ is a Legendre–Fenchel transform (by formula 5.4) and thus usc and concave on $W^{2,1}(X).$ Furthermore, $\mathcal{E}$ is also usc and concave on $W^{2,1}(X)$ (using the formula in Example 3.2 and the fact that the Hilbert norm on a Hilbert space is convex and continuous on the Hilbert space. Thus applying formula 5.5 concludes the proof of the duality formula 5.1.
@@ 5.1.2. Existence of a dual minimizer in $W^{2,1}(X)$
We next specialize to the log Fano curve $(X,\Delta)=(X,0)$ and take $\phi_{0}$ to be SU(2)SU(2)–invariant, so that $\mu_{0}$ and $\omega_{0}/2$ both coincide with $\omega_{\text{FS }}$ and hence with the $SO(3)$-invariant probability measure on $S^{2}.$ We will show that the functional $\mathcal{D}_{\gamma}(u)+\delta_{\gamma}(u)$ admits a minimizer in $W^{2,1}(X)$ (the same argument applies to any log Fano curve). To prove this we make a change of variable from $u$ to $v$ on $W^{2,1}(X),$ defined by
@@
Now take $\gamma^{\prime}>\gamma$ such that $\gamma^{\prime}<2.$ Combining Theorem 3.13 for $(X,0)$ with the duality formula 5.1 shows that there exists a constant $C$ such that
@@
The uniform $W^{2,1}$-bound 5.6 implies, that, after passing to a subsequence,
@@
This is standard, but for completeness we provide a proof. First, by the Rellich–Kondrachov theorem, the uniform bound 5.6 implies that, after passing to a subsequence, $v_{j}\rightarrow v_{\infty}$ in $L^{1}(X).$ As a consequence, after again passing to a subsequence, we may assume that $v_{j}(x)\rightarrow v_{\infty}(x)$ for almost any $x$ in $X.$ Next, by the Moser-Trudinger inequality on $X$ the uniform bound 5.6 also implies that, for any given $\delta>0,$ there exists $A_{\delta}>0$ such that
@@
using the uniform bound 5.8 for some $\delta>0.$ Finally, letting first $j\rightarrow\infty$ and using the dominated convergence theorem and then letting $M\rightarrow\infty$ and using the monotone convergence theorem, shows that $\int e^{2v_{j}}g\omega_{0}$ converges towards $\int e^{2v_{\infty}}g\omega_{0},$ as desired. ∎
All that remains is to show that $\delta(v)$ is lsc along the sequence $v_{j}.$ To this end, take $\mu_{j}\in\mathcal{P}(X)$ such that $D(\mu_{j}|e^{2v_{j}}\omega_{\text{FS}}/\int e^{2v_{j}}\omega_{\text{FS}})\leq\delta(v_{j})+1/j.$ Since $\mathcal{P}(X)$ is compact, we may, after passing to a subsequence, assume that $\mu_{j}\rightarrow\mu_{\infty}$ in $\mathcal{P}(X)$ for some $\mu_{\infty}.$ Since $D(\cdot|\cdot)$ is lsc with respect to the product topology on $\mathcal{P}(X)^{2}$ it thus follows from the weak convergence 5.7 that

### v3

@@ 4.1. Spontaneous symmetry breaking
although both $F_{-\gamma}$ and $\mathcal{P}(X)_{0}$ are $S^{1}$–invariant (and likewise for $\mathcal{D}_{\gamma}$ and $C^{1}(X)_{0})$. The point is that the destabilizing curve $\mu^{t}$, exhibited in the proof of Theorem 3.13, is not $S^{1}$–invariant (with respect to the $S^{1}$–action fixing 0 and $\infty).$ This is in contrast to the case when $\gamma=1,$ where the two infima above do coincide. Indeed, when $\gamma=-1$, the inf of $F_{-\gamma}$ is attained at the unique Kähler–Einstein volume form in $\mathcal{P}(X)_{0}$ and this volume form is $S^{1}$–invariant (and likewise for $\mathcal{D}_{1}(u)).$
@@ 4.1.1. Failure of symmetric rearrangements
In [43, 44] and [10], the sharp logarithmic HLS inequality on the two-sphere $X$ (corresponding to the case $w=0$ and $\gamma=1)$ was shown using symmetric decreasing rearrangements. The arrangement was defined with respect to Lebesgue measure on $\mathbb{R}^{2}$ in [43, 44] (using stereographic projection), while in [10] it was defined with respect to the $\mathrm{SO}(3)$–invariant measure on the two-sphere. Both these two rearrangements define maps $\mathcal{P}(X)\rightarrow\mathcal{P}(X)^{S^{1}}$ decreasing the functional $F_{-1}$ (and the arrangement in [10] decreases $F_{-\gamma}$ for any $\gamma>0).$ However, neither of the rearrangements preserve the space $\mathcal{P}(X)_{0}$ (for any given weight $w).$ Moreover, the spontaneous symmetry breaking exhibited in formula 4.12 implies that when $w\in]0,1/2[$ and $\gamma\in]1/(1-w),2[$ there is no map $\mathcal{P}(X)_{0}\rightarrow\mathcal{P}(X)_{0}^{S^{1}}$ that decreases the functional $F_{-\gamma}.$
@@ 4.2. Restricting to $\mathcal{K}$-invariant metrics
By [95, Thm 2.18], $\delta_{C(\mathcal{G})}^{A}(X,\Delta)>1$ iff $(X,\Delta)$ admits a KE metric. Hence, by [61, Thm 2.18], $\delta_{C(\mathcal{G})}^{A}(X,\Delta)>1$ iff $\delta_{\mathcal{G}}^{A}(X,\Delta)>1.$ However, in general these two analytic invariants may differ:
@@
5. A sharp constrained log HLS inequality on $S^{2}$ and application to quantitative stability
In this section we will prove the following result, which is equivalent to Theorem 1.6 stated in the introduction (by formulas 3.15, 3.11):
@@ 5. A sharp constrained log HLS inequality on $S^{2}$ and application to quantitative stability
Let $X$ be the two-sphere (viewed as the log pair $(X,0)$) and assume that $\omega_{0}\in c_{1}(-K_{X})$ is $\mathrm{SO}(3)$–invariant. Then for any given $\gamma\in]0,2]$
(5.1) $F_{-\gamma}(\mu):=-\gamma E(\mu)+D(\mu)\geq 0\,\,\,\forall\mu\in\mathcal{P}(X)_{0},$
where $\mathcal{P}(X)_{0}$ denotes the space of all probability measures on $X$ satisfying the moment condition $\int_{X}$$\boldsymbol{x}\mu=0$. Moreover, equality holds iff $\mu$ is the $\mathrm{SO}(3)$–invariant probability measure on $X.$ By Prop 3.4, this equivalently means that for any $\mathcal{\mu\in P}(X)$ such that $E(\mu)<\infty$
(5.2) $F_{-1}(\mu)\geq 1\cdot\inf E_{g\in\mathcal{G}}(g^{*}\mu)$
with equality iff $\mu$ is in the $\mathcal{G}$-orbit of $\mu_{0}.$
@@
We recall that, by Theorem 3.13, $F_{-\gamma}(\mu)$ is unbounded from below when $\gamma>2,$ which implies (by Prop 3.4 that the constant $1$ in the inequality 5.2 cannot be replaced by a strictly larger constant. When $\gamma=1$ the inequality 5.1 follows from the general fact that, when a Fano manifold $X$ admits a Kähler–Einstein metric $\omega_{\text{KE}}\in C^{1}(X),$ the Mabuchi functional $\mathcal{M}$ is minimized on the Kähler potential $u$ of $\omega_{\text{KE}}$ [9] and $\mathcal{M}$ is then $\mathcal{G}$-invariant – so that the moment condition on $\mu$ can be dispensed with. In the case of the two-sphere (or more generally, the unit sphere in $\mathbb{R}^{n+1})$ the inequality $F_{-1}(\mu)\geq 0$ on $\mathcal{P}(X)$ is shown in [43, 44, 10] using symmetrization (the inequality $F_{-\gamma}(\mu)\geq 0$ for $\gamma\in[0,1]$ then follows directly).
@@
Before embarking on the detailed proof of Theorem 5.1 it may be illuminating to compare its structure to the proof of Theorem 4.5 in [80]. The starting point of the latter is Aubin’s constrained Moser–Trudinger inequality which, in our notations, implies that the twisted Ding functional $\mathcal{D}_{\gamma}$ admits a minimizer $u_{\gamma}$ on the space of all functions in the Hilbert space $W^{2,1}(X)$ satisfying Aubin’s constraint. Since this is an infinite dimensional Hilbert manifold it follows that $u_{\gamma}$ satisfies an Euler–Lagrange equation, involving the differential of $\mathcal{D}_{\gamma}$ and three Lagrange multipliers (incorporating Aubin’s constraint). As shown in [48, Lemma 3.1], the Lagrange multipliers vanish and, as a result, $u_{\gamma}$ is, in fact, a critical point of $\mathcal{D}_{\gamma}$ on $W^{2,1}(X)$ and thus satisfies a Liouville equation on $S^{2},$ that is shown to have only have constant solutions in [80] (thus confirming a conjecture in [48]).
@@
In order to prove Theorem 5.1 we replace $\mathcal{D}_{\gamma}$ with the free energy functional $F_{-\gamma}$ on $\mathcal{P}(X)_{0}.$ It follows from the formula $\delta^{\mathrm{A}}(X,0)^{\mathcal{G}}=2$ in Theorem 3.13 that $F_{-\gamma}$ admits a minimizer $\mu_{\gamma}$ in $\mathcal{P}(X)_{0}$ for any given $\in]0,2[.$ However, since the space $\mathcal{P}(X)$ has a boundary it seems hard to directly show that $\mu_{\gamma}$ satisfies an appropriate equation. Instead, we leverage a duality formula, expressing the infimum of $F_{-\gamma}$ over $\mathcal{P}(X)_{0}$ as the infimum over $W^{2,1}(X)$ of a perturbation of the twisted Ding functional:
$\inf_{\mathcal{P}(X)_{0}}F_{-\gamma}=\gamma\inf_{u\in W^{2,1}(X)}\left(\mathcal{D}_{\gamma}(u)+\delta_{\gamma}(u)\right),$
@@
where $\delta_{\gamma}$ is an entropic distance, defining a function on $W^{2,1}(X)$ taking values in $0,\infty[,$ that vanishes precisely when $\gamma u$ satisfies Aubin’s constraint 4.9 (in particular, the unboundedness from above of $\delta_{\gamma}(u)$ compensates for the unboundedness from below of $\mathcal{D}_{\gamma}(u)).$ This bypasses the boundary problem and allows us to reduce the proof of Theorem 5.1 to the uniqueness result in [80]. In fact, the proof will show that the inf in the right hand side is attained precisely when $u$ is constant and thus vanishes.
@@
A consequence of the proof is that
(5.3) $\inf_{u\in W^{2,1}(X)}\left(\mathcal{D}_{\gamma}(u)+\delta_{\gamma}(u)\right)=\inf_{u\in W^{2,1}(X)}\left(\mathcal{D}_{\gamma}(u)+\chi(\gamma u)\right)$
where $\chi$ is the indicator function attached to Aubin’s constraint, i.e. $\chi(v)$ vanishes when $v$ satisfies Aubin’s constraint and $\chi(v)=\infty$ otherwise. From a general optimization perspective - where the infimum in the left hand side in equation 5.3 can be viewed as a relaxation of the constained infimum in the right hand side - this is somewhat miraculous. Indeed, a priori, the left hand side of is merely smaller then or equal to the right hand side, since $\delta_{\gamma}(u)\leq\chi(\gamma u)$ and the infimum in the left hand side could even equal -$\infty$ when the right hand side is finite. This is a reflection of the fact that Theorem 1.6 is, a priori, stronger than Theorem 4.5.
@@
which defines a probability measure on $X,$ by the previous lemma. It will be enough to prove the inequality in Theorem 5.1 for a fixed $\gamma\in]1,2[.$ By Theorem 3.13 the infimum of $F_{-\gamma}$ on $\mathcal{P}(X)_{0}$ is then finite. By the previous lemma we thus have
$\inf_{\mathcal{P}(X)_{0}}F_{-\gamma}=\inf_{\mathcal{P}(X)_{0}\cap W^{2,1*}}F_{-\gamma}>-\infty.$
@@ 5.1.1. A duality formula
(5.4) $\inf_{\mathcal{P}(X)_{0}\cap W^{2,1*}}F_{-\gamma}=\gamma\inf_{u\in W^{2,1}(X)}\left(\mathcal{D}_{\gamma}(u)+\delta_{\gamma}(u)\right),\,\,\,\delta_{\gamma}(u)\coloneqq\frac{1}{\gamma}\inf_{\mu\in\mathcal{P}(X)_{0}}D(\mu|\mu_{\gamma u}),$
@@
then proves 5.7.
Next, note that the functional $\tilde{D}(\mu)$ on $W^{2,1}(X)^{*}$ is convex and lsc (since $\mathcal{P}(X)_{0}$ is compact and convex in $W^{2,1}(X)^{*})$ and $D(\mu|\mu_{0})$ is convex on $\mathcal{P}(X)_{0}).$ It thus follows from combining formula 5.7 with the duality 5.5 that
@@
so that formula 5.6 translates into
$\inf_{\mathcal{P}(X)_{0}\cap W^{2,1}(X)^{*}}F_{-\gamma}=\inf_{W^{2,1}(X)^{*}}\left(-\mathcal{E}^{*}+\mathcal{A}^{*}\right).$
We will use the following essentially well-known identity:
@@
By 5.5, this means that $\mathcal{A}(u)\geq\mathcal{E}(u)+C$ for any $u\in H.$ Conversely, if the latter inequality holds, then, by taking the sup over all $u\in H$ we directly deduce that $\mathcal{A}^{*}(\mu)\geq\mathcal{E^{*}}(\mu)+C$ for all $\mu\in H^{*}.$ Thus $(-\mathcal{E}^{*}+\mathcal{A}^{*})(\mu)\geq C$ for all $\mu\in H^{*}.$
In the present case $\mathcal{A}$ is a Legendre–Fenchel transform (by formula 5.7) and thus usc and concave on $W^{2,1}(X).$ Furthermore, $\mathcal{E}$ is also usc and concave on $W^{2,1}(X)$ (using the formula in Example 3.2 and the fact that the Hilbert norm on a Hilbert space is convex and continuous on the Hilbert space. Thus applying formula 5.8 concludes the proof of the duality formula 5.4.
@@ 5.1.2. Existence of a dual minimizer in $W^{2,1}(X)$
We next specialize to the log Fano curve $(X,\Delta)=(X,0)$ and take $\phi_{0}$ to be $\mathrm{SU}(2)$–invariant, so that $\mu_{0}$ and $\omega_{0}/2$ both coincide with $\omega_{\text{FS }}$ and hence with the $\mathrm{SO}(3)$-invariant probability measure on $S^{2}.$ We will show that the functional $\mathcal{D}_{\gamma}(u)+\delta_{\gamma}(u)$ admits a minimizer in $W^{2,1}(X)$ (the same argument applies to any log Fano curve). To prove this we make a change of variable from $u$ to $v$ on $W^{2,1}(X),$ defined by
@@
Now take $\gamma^{\prime}>\gamma$ such that $\gamma^{\prime}<2.$ Combining Theorem 3.13 for $(X,0)$ with the duality formula 5.4 shows that there exists a constant $C$ such that
@@
The uniform $W^{2,1}$-bound 5.9 implies, that, after passing to a subsequence,
@@
This is standard, but for completeness we provide a proof. First, by the Rellich–Kondrachov theorem, the uniform bound 5.9 implies that, after passing to a subsequence, $v_{j}\rightarrow v_{\infty}$ in $L^{1}(X).$ As a consequence, after again passing to a subsequence, we may assume that $v_{j}(x)\rightarrow v_{\infty}(x)$ for almost any $x$ in $X.$ Next, by the Moser-Trudinger inequality on $X$ the uniform bound 5.9 also implies that, for any given $\delta>0,$ there exists $A_{\delta}>0$ such that
@@
using the uniform bound 5.11 for some $\delta>0.$ Finally, letting first $j\rightarrow\infty$ and using the dominated convergence theorem and then letting $M\rightarrow\infty$ and using the monotone convergence theorem, shows that $\int e^{2v_{j}}g\omega_{0}$ converges towards $\int e^{2v_{\infty}}g\omega_{0},$ as desired. ∎
All that remains is to show that $\delta(v)$ is lsc along the sequence $v_{j}.$ To this end, take $\mu_{j}\in\mathcal{P}(X)$ such that $D(\mu_{j}|e^{2v_{j}}\omega_{\text{FS}}/\int e^{2v_{j}}\omega_{\text{FS}})\leq\delta(v_{j})+1/j.$ Since $\mathcal{P}(X)$ is compact, we may, after passing to a subsequence, assume that $\mu_{j}\rightarrow\mu_{\infty}$ in $\mathcal{P}(X)$ for some $\mu_{\infty}.$ Since $D(\cdot|\cdot)$ is lsc with respect to the product topology on $\mathcal{P}(X)^{2}$ it thus follows from the weak convergence 5.10 that